### v1

# Does synthetic data help low-resource Named Entity Recognition?

Abstract
Named Entity Recognition(NER) for low-resource languages aims to produce robust systems for languages where there is limited labeled training data available, and has been an area of increasing interest within NLP. Data augmentation for increasing the amount of low-resource labeled data is a common practice. In this paper, we explore the role of synthetic data in the context of multilingual, low-resource NER, considering 11 languages from diverse language families. Our results suggest that synthetic data does in fact hold promise for low-resource language NER, though we see significant variation between languages.
Does synthetic data help low-resource Named Entity Recognition?
Gaurav Kamaththanks: work done during an internship at the National Research Council, Canada McGill University, Canada gaurav.kamath@mail.mcgill.ca                        Sowmya Vajjala National Research Council, Canada sowmya.vajjala@nrc-cnrc.gc.ca
1 Introduction
As a sequence labeling task, named entity recognition (NER) has historically one of the well-studied tasks in NLP. While much of the research in this direction assumes the presence of large datasets Sang and De Meulder (2003); Weischedel et al. (2011), there is also a large body of work focused on low-resource NER, i.e., developing methods for data scarce scenarios. Such scenarios can exist in two contexts: low-resource languages (e.g., Sohn et al., 2024) as well as low-resource domains within English (e.g., Santoso et al., 2024), both of which received considerable attention in NLP research in recent years. We approach the problem in the low-resource language context in this paper.
Figure 1: High-level overview of our data generation process. We use multilingual large language models to generate new NER datapoints on the basis of a handful of high quality datapoints. See Section 3.4 for more.
Two common approaches to address such data scarce scenarios are: knowledge transfer and data augmentation. While knowledge transfer involves using cross-linguistically compatible text representations and/or transfer learning from a related high-resource setup, data augmentation approaches focus on increasing the amount of labeled data in the target setup. There are also large automatically constructed multilingual NER datasets that rely on large multilingual data sources such as Wikipedia Pan et al. (2017); Malmasi et al. (2022) and some such datasets have become a part of large multilingual benchmarks Asai et al. (2024). Recent research explored the usefulness of Large Language Models for such data augmentation and synthetic data creation in NLP Whitehouse et al. (2023); Li et al. (2023), and their use for NER is emerging Bogdanov et al. (2024).
In this background, we propose LLM based synthetic data generation using a small amount of gold examples as an alternative to relying on large scale, automatically labeled datasets for low-resource NER. By considering 11 languages from 3 language families- Danish, Swedish and Slovak from the Indo-European languages; Swahili, Kinyarwanda, Yoruba and Igbo from the Niger-Congo languages, and Kannada, Malayalam, Tamil, and Telugu from the Dravidian languages, we show that:
  1. 1.
    Creating synthetic data by prompting a large language model with a few manually labeled examples gives much better results than completely automatically labeled datasets.
  2. 2.
    Zero-shot transfer after training for NER on a related language is a surprisingly high baseline.
  3. 3.
    A small amount of high quality, manually annotated training data is often better than a large amount of synthetic data in low-resource NER performance.
To our knowledge, this is the first attempt to explore using large language models for synthetic data generation in low-resource NER, and the first to cover $>10$ languages. Based on our findings, we also summarize some best-practices for synthetic data generation with LLMs for low-resource NER.
We start with a review of related literature (Section 2) and describe our data generation approach in Section 3. The Results are detailed in Section 4 followed by further discussion on limitations and broader impact.
2 Related Work
Addressing NER for low resource scenarios has been one of the topics of interest in NLP for a long time. We can broadly identify three strands of research on this issue: One strand of research focused on how to achieve good transfer from a source language to the target language with a much smaller amount of training data. With the arrival of pre-trained language models, this extended into mono/multilingual NER models that could be used for zero/few-shot transfer on low(er) resource languages. Another direction of research focused on creating large scale datasets to address the data scarcity issue in low-resource languages. We summarize low-resource NER research along these two strands in this section.
Cross-lingual Transfer in NER:
Some of the early approaches in cross-lingual transfer relied on the use of dictionaries, knowledge bases, and parallel data to enrich the feature representations for NER in the target language Burkett et al. (2010); Darwish (2013). The use of non-lexical features such as POS tags, multilingual gazetteers Zirikly and Hagiwara (2015), phonological features Bharadwaj et al. (2016), translation and transliteration Huang et al. (2019); Johnson et al. (2019), and learning joint character representations between related languages Cotterell and Duh (2017) are some of the ideas explored in this context. Research advances in multilingual word and text embeddings, followed by pre-trained models, lead towards exploring the possibility of zero-shot or few-shot NER through a source NER model trained using labeled data from one or multiple languages Rahimi et al. (2019); Mueller et al. (2020); Zeng et al. (2022); Zhao et al. (2022); Yang et al. (2022); Zhou et al. (2022).
Creating Synthetic Datasets for Low-resource NER:
Parallel corpora or machine translated versions, and multilingual gazetteers were used as a means to create low-resource NER datasets from scratch by utilizing high quality NER datasets from a high resource language in past research Mayhew et al. (2017); Ni et al. (2017); Xie et al. (2018); Liu et al. (2021); Yang et al. (2022); Fetahu et al. (2022). WikiAnn Pan et al. (2017) is a large scale NER dataset for 282 languages, created by mining named entity data from Wikipedia’s cross-lingual links and has been a popular benchmark for cross-lingual and multilingual NER over the past years. However, some recent work Lignos et al. (2022) pointed to the low-quality example sets, and the lack of a human verified test set. Despite this, the dataset continues to be a common benchmark for cross-lingual NER Schmidt et al. (2022); Asai et al. (2024).
LLMs as Data Generators:
Some of the recent work explored using large language models such as GPT3.5 as data generators for NER. NuNER Bogdanov et al. (2024) uses GPT3.5 to automatically annotate a subset of the English part of C4, a cleaned common crawl dataset Raffel et al. (2020) and demonstrated its extendability to 7 other Indo-European languages. ProgGen Heng et al. (2024) proposed an approach to use GPT4 to generate domain relevant attributes which are then utilized to create domain specific training data and evaluated their approach on several English NER datasets across domains. Wang (2024).
Compared to these two approaches, we study synthetic data generation for NER across 11 languages covering three language families and five scripts, by taking high-quality labeled sentences as seed examples to generate new data, instead of relying on completely automatically created datasets. We also evaluate open LLMs along with GPT4 with this approach. Finally, we compare this approach with WikiAnn Pan et al. (2017); Rahimi et al. (2019), and with cross-lingual transfer from a related language with high-quality data.
3 Our Approach
We describe our general experimental setup for training and fine-tuning of NER models and the synthetic data generation process in this section.
3.1 Experimental Setup
At a high level, our approach to assess the potential of synthetic data for low-resource language NER involves the following steps:
  1. 1.
    Using the train split of some high quality, usually manually annotated NER dataset for a target language to generate synthetic data for that language with the help of an LLM, and
  2. 2.
    Comparing the performance of an NER model on the test split of the high quality dataset when trained on synthetic data from Step 1 and another model trained on the train split of the high quality dataset.
For all of our experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model. We fine-tune this model on our synthetic and original training sets in two distinct settings.
  1. 1.
    In the first setting, we use our data to train an NER model from scratch, by fine-tuning the pre-trained XLM-RoBERTa-large on target language NER data.
  2. 2.
    In the second setting, we first fine-tune the model on the high quality NER data in a language related to the target language [^1] See Appendix for the full list of chosen related languages for all the target languages., so as to get a working NER model in the related language, and then further fine-tune this NER model on our synthetic or organic target language data.
While the first setting—which we name ner from scratch—aims to shed light on the relative utility of synthetic data for training an NER model (largely) from the ground up, the latter —which we name ner fine-tuning—simulates a common setting, when a lower resource language lacks adequate NER data, but is related to a higher-resource language with existing NER systems. In both settings, we modulate the amount of data (both synthetic and organic) used, so as to compare model performance when trained on smaller or larger amounts of each type of data.
3.2 Implementation
We source the pre-trained XLM-RoBERTa-large weights from Huggingface using the transformers library; fine-tuning is implemented using training pipelines from the same library. In the ner from scratch setting, we train on the the target language data for 10 epochs; in the ner fine-tuning setting, we train on the related language data for 5 epochs, and then the target language data for 10 epochs. In all cases, we use a learning rate of 2e-05, and a batch size of 16.
3.3 Languages & Datasets
LLM Response Quality Examples
Well-Formatted {"data": [{ "tokens": ["Lars", "Løkke", "Rasmussen", "besøgte", "firmaet", "i", "Odense", "."],
"ner_tags": [1, 2, 2, 0, 0, 0, 5, 0]}, …
Unequal Token & Tag Lengths {"id": "4123", "tokens": ["Wananchi", "wamekunja", "mashitaka", "."],
"ner_tags": [0, 0, 0, 0, 0, 0]}
Run-On & Incomplete Data {"id": "9000", "tokens": ["Olorun", "lẹ̀ẹ̀", ..., "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò", "ò"
{"id":"4617","tokens":["Ọḍ́ǹ", "ọ̀sè", "lè", "ọ̀rị́", "-", "èdè", "Ọḅ́f̣́emi", "àfú", "fùn", "àwọn", "gb", "." ]
"ner_tags":[8,0,0,0,0
Empty Responses & Prompt Continuations <EOS_TOKEN>
<EOS_TOKEN>include a mix of names, locations, organizations...
Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested.
We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European). Of these, Igbo, Yoruba, and Kinyarwanda are not a part of the 100 languages covered in the pre-trained XLM-Roberta model, but Swahili is a part of the pre-training languages.
We use the Universal NER dataset Mayhew et al. (2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al. (2022) for Kinyarwanda, Swahili, Igbo and Yoruba, and the Naamapadam dataset Mhaske et al. (2023) for Tamil, Kannada, Malayalam and Telugu. It is worth noting that while the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus, in the authors’ admission, contain some amount of noise. Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them. The Universal NER and Naamapadam datasets cover persons, locations and organizations as categories; the MasakhaNER2 data covers these three categories, as well as dates.
3.4 Synthetic Data Generation Process
Our synthetic data generation process involves using large language models to generate new synthetic datapoints on the basis of existing, high quality NER annotations through the following steps:
  1. 1.
    First, we randomly sample $m$ data points from the train split of an organic NER dataset.
  2. 2.
    Next, we format and append these data points to a prompt asking the model to produce $n$ new, unique data points on the basis of the $m$ datapoints in the prompt.
  3. 3.
    We submit this prompt as input to the LLM, and extract any correctly-formatted data points from its response;
  4. 4.
    We repeat steps (1)-(3) $k$ times, with each call to the model choosing a different random sample of organic data points.
Figure 1 shows a demonstration of this approach, with the exact prompt used. In our experiments, we set $m$ to 10, $n$ to 20, and $k$ to 500. This sets an upper cap of 5000 synthetic training data points, if every model response contains perfectly formatted data points.
Our method to extract correctly-formatted datapoints from raw model responses relies on regular expressions to identify potential JSON substrings in the model response; as shown in Table 1, model responses may contain data that is ill-formatted due to a number of reasons. We deem any data faulty for such reasons (among any other reasons, such as hallucinated labels) to be unusable, and do not include it in our synthetic train sets.
We experiment with three LLMs as our source of synthetic data. We use GPT-4222We use gpt-4-turbo, and all data generation with the model was conducted between November and December 2024. Achiam et al. (2023), which we assume to be the state of the art; we use Llama-3.1-8B-Instruct Dubey et al. (2024), as an instance of a much smaller, open-source instruction-tuned model; and finally, we use Cohere For AI’s aya-expanse-32b Dang et al. (2024), as an instance of a larger open source multilingual large language.
4 Results
4.1 Qualitative Analysis of LLM-Generated Data
Model Average % of Usable Train Datapoints
(Across All Languages)
GPT-4-turbo 97.0%
Llama-3.1-8B-Instruct* 59.3%
aya-expanse-32b 11.7%
Table 2: Average percentage of usable train datapoints produced by each model; we divide the number of usable synthetic datapoints produced for each language by 5000 (the synthetic dataset size if all responses were perfect), and then take the average across languages. *Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits.
(a) ner from scratch Setting
(b) ner fine-tuning Setting
Figure 2: NER model performance on a subset of languages when trained on increasingly large subsets of organic and synthetic data. aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as the organic or GPT-4-produced data in fine-tuning data size. Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model. See Appendix for similar plots of all 11 languages.
Table 1 illustrates the type of responses we get from the three models when we prompt them for synthetic data in the manner described in Section 3.4.
Broadly speaking, we find that GPT-4 provides the highest quality datapoints, followed by Llama-3.1-8B-Instruct, and lastly aya-expanse-32b. This is equally borne out by the proportion of usable training datapoints produced by each model, shown in Table 2.
While model responses from GPT-4 are almost always usable, we find a few persistent errors in responses from the other two models. We highlight some below:
  1. 1.
    Unequal token and tag lengths: Some responses from a model are largely well-formatted and coherent, but are rendered unusable because of a mismatch in their numbers of tokens and tags.
  2. 2.
    Run-on responses and incomplete data: Some responses from the models result in endless repeated sequences; in other cases, responses hit new token limits before arriving at complete datapoints.
  3. 3.
    Empty responses and prompt continuations: These were specific to aya-expanse-32b; model responses to the generation prompt often either terminated with <EOS> tokens, or simply continued the generation prompt, as shown in the example in Table 1.
Given the wide range of languages we test, and the cost and difficulty of finding trained annotators to manually annotate samples from all 11 languages, we do not source human annotations of the synthetic data to compare against. Instead, as detailed in Section 3.1, we assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
4.2 Training on Synthetic Data
Figure 2 shows our results on a subset of languages—Danish, Tamil and Igbo—when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings (see Appendix for such plots across all languages). As the figure illustrates, in all cases, models trained on LLM-generated data struggle to compete with those trained on the original, organic train sets. Perhaps more surprisingly, in the ner fine-tuning setting, zero-shot transfer onto the target language outperforms the same models after they have been further fine-tuned on synthetic target language data, suggesting that in such cases, the synthetic data is actually detrimental to their performance.
In terms of comparisons between data-generating LLMs, these results also confirm our initial qualitative observations (see Section 4.1). Models trained on synthetic data from GPT-4 do better than those trained on synthetic data from Llama-3.1-8B-Instruct, which in turn outperform the aya-expanse-32b equivalents—suggesting that the quality of data produced by the three LLMs follows the same order.
4.3 WikiAnn
Language n.f.s. F1 n.f.t. F1 Data Size
Kannada WikiANN 4.5e-3 0.77 100
GPT-4 0.26 0.65 100
GPT-4 0.65 0.68 4861
Naamapadam 0.47 0.79 100
Naamapadam 0.76 0.79 5000
Danish WikiANN 0.72 0.71 20000
GPT-4 0.60 0.68 4857
Universal NER 0.83 0.85 4383
Yoruba WikiANN 0.07 0.21 100
GPT-4 0.26 0.43 100
GPT-4 0.53 0.56 4761
MasakhaNER 2 0.20 0.50 100
MasakhaNER 2 0.79 0.82 5000
Swahili WikiANN 0.50 0.59 1000
GPT-4 0.74 0.78 1000
GPT-4 0.75 0.79 4900
MasakhaNER 2 0.69 0.85 1000
MasakhaNER 2 0.92 0.90 5000
Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for a handful of languages. See Appendix for full table of results. n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.
As a further evaluation of the viability of synthetically produced data for lower-resource language NER, we also extend our comparison to the WikiANN dataset. The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al. (2017); Rahimi et al. (2019) [^3] As Lignos et al. (2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al. (2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al. (2019), contains data from 176 languages. In this work, we refer to the latter when referring to the WikiANN dataset.. Though it is often used as a multilingual NER benchmark (e.g. Asai et al., 2024), however, the WikiANN dataset in fact does not include manual annotations, and is instead produced by scraping Wikipedia data. Consequently, the dataset represents a different sort of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases.
Table 3 shows the performance of NER models trained on WikiANN data, as compared with those trained on our organic and LLM-produced synthetic data. The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g. Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints). In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
As Table 3 indicates, training on WikiANN data yields NER models that perform considerably worse than those trained on organic data, and often worse than those trained on synthetic data produced by GPT-4 (when data size is equal between the two). The WikiANN data for some languages appears completely ill-suited for training an NER model in the ner from scratch setting, while, perhaps unsurprisingly, organic data produces the best results.
5 Conclusions and Discussion
Our results suggest three broad findings around the utility of LLM-generated synthetic data for low resource language NER.
  1. 1.
    The first, and perhaps most significant, is that a small amount of carefully annotated data yields better performance than a large amount of synthetic data. As is evident in Figure 2, even 100 manually annotated datapoints can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
  2. 2.
    The second is that, in many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may actually lower its performance.We find this to be true in the case of all languages tested except the four Niger-Congo languages—Igbo, Yoruba, Swahili and Kinyarwanda. In these cases, it is worth noting that the overall baselines are lower, presumably because these languages are all lower resource than the others tested.This may explain why synthetic data yields performance gains over the zero-shot baseline, though it does not change the trend of a small amount of manually annotated data yielding far better performance.
  3. 3.
    Lastly, we find that LLM-generated synthetic data often produces better results than automatically labelled data. Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN data, as shown in the case of Kannada, Yoruba and Swahili. This echoes the findings by Lignos et al. (2022), who arrive at similarly negative findings around the data quality of WikiANN; consequently, they further reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
6 Limitations
Although we experimented with many languages, the nature of the NER datasets used is relatively simple. Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories. While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data. Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation. So, our findings should be viewed after taking these aspects into consideration.
7 Ethics and Broader Impact
We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work. The models built using such artificially created datasets should always be validated with a human-labeled data. We did not involve any human participants in this study. All the code and generated datasets will be provided as supplementary material to support reproducible research.
Acknowledgments
This research was conducted at the National Research Council of Canada, thereby establishing a copyright belonging to the Crown in Right of Canada, that is, to the Government of Canada.
Appendix A Related-Language Model Details
In the ner fine-tuning setting, we first train an NER model on a language related to the target language, before fine-tuning it further on the target language NER data. Below is the list of related languages chosen to build a base NER model for each target language.
Target Language Related Language Chosen
Kannada Telugu
Tamil Telugu
Telugu Kannada
Malayalam Tamil
Kinyarwanda Swahili
Swahili Kinyarwanda
Yoruba Igbo
Igbo Yoruba
Swedish Danish
Danish Swedish
Slovak English*
Table 4: List of related languages used in the ner fine-tuning setting for each target language. *English is not closely related to Slovak, but given the absence of another highly related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
Appendix B LLM Prompts & Implementation
As shown in Figure 1, we present the following prompt to the LLM in the data generation process:
Help me make a {language} Named Entity Recognition dataset. Please give me {n} new datapoints, formatted as a single JSON object. Make sure the examples are unique and diverse. Here are some examples to get you started:
{m examples}
For GPT-4, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings. For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt. For the open-sourced models, we used the following system prompt:
You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce. You only output JSON strings.
For GPT-4, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce.
We ran all open-sourced models using vLLM Kwon et al. (2023), with a temperature setting of 0.8, maximum new token limit of 4096 new tokens, and nucleus sampling value of 0.8. Calls to GPT-4 were made using default hyperparameters.
Appendix C Full Results of Training on Synthetic Data
C.1 NER From Scratch Setting
Figure 3 is a plot of results of training on synthetic data, in the ner from scratch setting, for all languages.
C.2 NER Fine-tuning Setting
Figure 4 is a plot of results of training on synthetic data, in the ner fine-tuning setting, for all languages.
Appendix D Full Results of WikiANN Comparison
Below in Table 5 is the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data.
Language n.f.s. F1 n.f.t. F1 Data Size
Kannada WikiANN 4.5e-3 0.77 100
GPT-4 0.26 0.65 100
GPT-4 0.65 0.68 4861
Naamapadam 0.47 0.79 100
Naamapadam 0.76 0.79 5000
Telugu WikiANN 0.67 0.74 1000
GPT-4 0.15 0.68 100
GPT-4 0.67 0.72 4919
Naamapadam 0.50 0.79 100
Naamapadam 0.83 0.82 5000
Tamil WikiANN 0.55 0.62 15000
GPT-4 0.56 0.51 4977
Naamapadam 0.73 0.73 5000
Malayalam WikiANN 0.65 0.74 10000
GPT-4 0.64 0.70 4898
Naamapadam 0.79 0.83 5000
Yoruba WikiANN 0.07 0.21 100
GPT-4 0.26 0.43 100
GPT-4 0.53 0.56 4761
MasakhaNER 2 0.20 0.50 100
MasakhaNER 2 0.79 0.82 5000
Swahili WikiANN 0.50 0.59 1000
GPT-4 0.74 0.78 1000
GPT-4 0.75 0.79 4900
MasakhaNER 2 0.69 0.85 1000
MasakhaNER 2 0.92 0.90 5000
Kinyarwanda WikiANN 7.9e-4 0.35 100
GPT-4 0.23 0.46 100
GPT-4 0.58 0.54 4754
MasakhaNER 2 0.26 0.61 100
MasakhaNER 2 0.80 0.81 5000
Igbo WikiANN 7.7e-3 0.39 100
GPT-4 0.43 0.70 100
GPT-4 0.66 0.71 4693
MasakhaNER 2 0.41 0.72 100
MasakhaNER 2 0.81 0.86 5000
Danish WikiANN 0.72 0.71 20000
GPT-4 0.60 0.68 4857
Universal NER 0.83 0.85 4383
Swedish WikiANN 0.36 0.29 20000
GPT-4 0.65 0.56 4825
Universal NER 0.58 0.89 4303
Slovak WikiANN 0.57 0.55 20000
GPT-4 0.29 0.29 4889
Universal NER 0.80 0.82 5000
Table 5: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for all 11 languages. n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.
Figure 3: ner from scratch Setting
Figure 4: ner fine-tuning Setting

### v2

@@
-Does synthetic data help low-resource Named Entity Recognition?
+Does Synthetic Data Help Named Entity Recognition for Low-Resource Languages?
 Abstract
-Named Entity Recognition(NER) for low-resource languages aims to produce robust systems for languages where there is limited labeled training data available, and has been an area of increasing interest within NLP.
-Data augmentation for increasing the amount of low-resource labeled data is a common practice.
-In this paper, we explore the role of synthetic data in the context of multilingual, low-resource NER, considering 11 languages from diverse language families.
-Our results suggest that synthetic data does in fact hold promise for low-resource language NER, though we see significant variation between languages.
-Does synthetic data help low-resource Named Entity Recognition?
-Gaurav Kamaththanks: work done during an internship at the National Research Council, Canada McGill University, Canada gaurav.kamath@mail.mcgill.ca                        Sowmya Vajjala National Research Council, Canada sowmya.vajjala@nrc-cnrc.gc.ca
+In this paper, we explore whether synthetic datasets generated by large language models are useful for low-resource named entity recognition, considering 11 languages from diverse language families.
+Our results suggest that synthetic data created with seed human labeled data is a reasonable choice when there is no available labeled data, and is better than using automatically labeled data.
+However, a small amount of high-quality data, coupled with cross-lingual transfer from a related language, always offers better performance.
+Does Synthetic Data Help
+Named Entity Recognition for Low-Resource Languages?
+Gaurav Kamaththanks: Work done during an internship at the National Research Council, Canada.1, 2  Sowmya Vajjala3 1McGill University, Canada  2Mila - Quebec AI Institute, Canada 3National Research Council, Canada gaurav.kamath@mail.mcgill.casowmya.vajjala@nrc-cnrc.gc.ca
 1 Introduction
-As a sequence labeling task, named entity recognition (NER) has historically one of the well-studied tasks in NLP.
-While much of the research in this direction assumes the presence of large datasets Sang and De Meulder (2003); Weischedel et al.
-(2011), there is also a large body of work focused on low-resource NER, i.e., developing methods for data scarce scenarios.
-Such scenarios can exist in two contexts: low-resource languages (e.g., Sohn et al., 2024) as well as low-resource domains within English (e.g., Santoso et al., 2024), both of which received considerable attention in NLP research in recent years.
-We approach the problem in the low-resource language context in this paper.
+Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP).
+Two common approaches to address this data scarcity are cross-lingual transfer and data augmentation/synthesis; recent research has in particular explored the usefulness of large language models (LLMs) for such data augmentation and synthetic data creation in NLP Whitehouse et al.
+(2023); Li et al.
+(2023), while their use for NER is also emerging Bogdanov et al.
+(2024).
+In this background, we propose LLM-based synthetic data generation using a small amount of gold examples (Figure 1) as an alternative to relying on automatically created datasets for low-resource NER.
+With experiments covering 11 languages, we show that
+  1. 1.
+    Even a small amount of human annotated data can yield far better performance than much larger amounts of synthetic data.
+  2. 2.
+    Zero-shot transfer from a related language can provide high baselines for low-resource language NER.
+  3. 3.
+    Synthetic data generated by prompting an LLM with a few high quality examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
 Figure 1: High-level overview of our data generation process.
@@
 See Section 3.1 for more.
-Two common approaches to address such data scarce scenarios are: knowledge transfer and data augmentation.
-While knowledge transfer involves using cross-linguistically compatible text representations and/or transfer learning from a related high-resource setup, data augmentation approaches focus on increasing the amount of labeled data in the target setup.
-There are also large automatically constructed multilingual NER datasets that rely on large multilingual data sources such as Wikipedia Pan et al.
-(2017); Malmasi et al.
-(2022) and some such datasets have become a part of large multilingual benchmarks Asai et al.
-(2024).
-Recent research explored the usefulness of Large Language Models for such data augmentation and synthetic data creation in NLP Whitehouse et al.
-(2023); Li et al.
-(2023), and their use for NER is emerging Bogdanov et al.
-(2024).
-In this background, we propose LLM based synthetic data generation using a small amount of gold examples as an alternative to relying on large scale, automatically labeled datasets for low-resource NER.
-By considering 11 languages from 3 language families- Danish, Swedish and Slovak from the Indo-European languages; Swahili, Kinyarwanda, Yoruba and Igbo from the Niger-Congo languages, and Kannada, Malayalam, Tamil, and Telugu from the Dravidian languages, we show that:
-  1. 1.
-    Creating synthetic data by prompting a large language model with a few manually labeled examples gives much better results than completely automatically labeled datasets.
-  2. 2.
-    Zero-shot transfer after training for NER on a related language is a surprisingly high baseline.
-  3. 3.
-    A small amount of high quality, manually annotated training data is often better than a large amount of synthetic data in low-resource NER performance.
-To our knowledge, this is the first attempt to explore using large language models for synthetic data generation in low-resource NER, and the first to cover $>10$ languages.
-Based on our findings, we also summarize some best-practices for synthetic data generation with LLMs for low-resource NER.
-We start with a review of related literature (Section 2) and describe our data generation approach in Section 3.
-The Results are detailed in Section 4 followed by further discussion on limitations and broader impact.
+We start with a review of related literature (Section 2) and describe our data generation approach and experimental setup in Section 3, followed by a discussion of the results (Section 4), limitations (Section 6) and broader impact (Section 7).
 2 Related Work
-Addressing NER for low resource scenarios has been one of the topics of interest in NLP for a long time.
-We can broadly identify three strands of research on this issue: One strand of research focused on how to achieve good transfer from a source language to the target language with a much smaller amount of training data.
-With the arrival of pre-trained language models, this extended into mono/multilingual NER models that could be used for zero/few-shot transfer on low(er) resource languages.
-Another direction of research focused on creating large scale datasets to address the data scarcity issue in low-resource languages.
-We summarize low-resource NER research along these two strands in this section.
-Cross-lingual Transfer in NER:
-Some of the early approaches in cross-lingual transfer relied on the use of dictionaries, knowledge bases, and parallel data to enrich the feature representations for NER in the target language Burkett et al.
-(2010); Darwish (2013).
-The use of non-lexical features such as POS tags, multilingual gazetteers Zirikly and Hagiwara (2015), phonological features Bharadwaj et al.
-(2016), translation and transliteration Huang et al.
-(2019); Johnson et al.
-(2019), and learning joint character representations between related languages Cotterell and Duh (2017) are some of the ideas explored in this context.
-Research advances in multilingual word and text embeddings, followed by pre-trained models, lead towards exploring the possibility of zero-shot or few-shot NER through a source NER model trained using labeled data from one or multiple languages Rahimi et al.
+NER in low resource settings has long been a topic of interest in NLP.
+Significant research examines cross-lingual transfer from a high resource source language to a lower-resource target language for the task Rahimi et al.
 (2019); Mueller et al.
@@
 (2022); Zhou et al.
-(2022).
-Creating Synthetic Datasets for Low-resource NER:
-Parallel corpora or machine translated versions, and multilingual gazetteers were used as a means to create low-resource NER datasets from scratch by utilizing high quality NER datasets from a high resource language in past research Mayhew et al.
+(2022), while other approaches have explored the creation of synthetic datasets through e.g.
+parallel corpora or machine translation Mayhew et al.
 (2017); Ni et al.
+(2017); Pan et al.
 (2017); Xie et al.
@@
 (2022).
-WikiAnn Pan et al.
-(2017) is a large scale NER dataset for 282 languages, created by mining named entity data from Wikipedia’s cross-lingual links and has been a popular benchmark for cross-lingual and multilingual NER over the past years.
-However, some recent work Lignos et al.
-(2022) pointed to the low-quality example sets, and the lack of a human verified test set.
-Despite this, the dataset continues to be a common benchmark for cross-lingual NER Schmidt et al.
-(2022); Asai et al.
+More recent work has explored using LLMs such as GPT-3.5 and GPT-4 as data generators for NER Bogdanov et al.
+(2024); Heng et al.
 (2024).
-LLMs as Data Generators:
-Some of the recent work explored using large language models such as GPT3.5 as data generators for NER.
-NuNER Bogdanov et al.
-(2024) uses GPT3.5 to automatically annotate a subset of the English part of C4, a cleaned common crawl dataset Raffel et al.
-(2020) and demonstrated its extendability to 7 other Indo-European languages.
-ProgGen Heng et al.
-(2024) proposed an approach to use GPT4 to generate domain relevant attributes which are then utilized to create domain specific training data and evaluated their approach on several English NER datasets across domains.
-Wang (2024).
-Compared to these two approaches, we study synthetic data generation for NER across 11 languages covering three language families and five scripts, by taking high-quality labeled sentences as seed examples to generate new data, instead of relying on completely automatically created datasets.
-We also evaluate open LLMs along with GPT4 with this approach.
-Finally, we compare this approach with WikiAnn Pan et al.
-(2017); Rahimi et al.
-(2019), and with cross-lingual transfer from a related language with high-quality data.
+We build on such work, but differ from their methods.
+Our data generation process uses high quality examples as seeds, and we not only evaluate different LLMs (both open and closed-source), but also experiment with 11 languages covering three language families and five base scripts.
 3 Our Approach
-We describe our general experimental setup for training and fine-tuning of NER models and the synthetic data generation process in this section.
-3.1 Experimental Setup
-At a high level, our approach to assess the potential of synthetic data for low-resource language NER involves the following steps:
+At a high level, our approach involves two steps:
   1. 1.
-    Using the train split of some high quality, usually manually annotated NER dataset for a target language to generate synthetic data for that language with the help of an LLM, and
+    Using the train split of a high quality (usually manually annotated) NER dataset for a target language to generate synthetic data for that language with the help of an LLM (Section 3.1); and then
   2. 2.
-    Comparing the performance of an NER model on the test split of the high quality dataset when trained on synthetic data from Step 1 and another model trained on the train split of the high quality dataset.
-For all of our experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al.
-(2020) as our base model.
-We fine-tune this model on our synthetic and original training sets in two distinct settings.
+    Comparing the performance of an NER model on the test split of the high quality dataset when trained on synthetic data from Step 1 and another model trained on the train split of the same high quality dataset (Section 3.2).
+3.1 Synthetic Data Generation:
+Our synthetic data generation process (shown in Figure 1) involves using LLMs to generate new synthetic data points on the basis of existing, high quality NER annotations as described below:
+  • First, we randomly sample $m$ data points from the train split of an organic (i.e. non-synthetic) NER dataset.
+  • Next, we format and append these data points to a prompt asking the model to produce $n$ new, unique data points on the basis of the $m$ data points in the prompt.
+  • We submit this prompt as input to the LLM, and extract any correctly-formatted data points from its response;
+  • We repeat steps (1)-(3) $k$ times, with each call to the model choosing a different random sample of organic data points.
+In our experiments, we set $m$ to 10, $n$ to 20, and $k$ to 500.
+This sets an upper cap of 5000 synthetic training data points, if every model response contains perfectly formatted data points.
+We present and solicit data structured as JSON strings to the LLMs, and extract well-formatted samples from model responses using regular expressions.
+Appendix A provides further details about this process.
+We compare three LLMs as our source of synthetic data: GPT-4111We use gpt-4-turbo, and all data generation with the model was conducted between September and December 2024.
+Achiam et al.
+(2023), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al.
+(2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al.
+(2024), as a larger open source multilingual LLM.
+3.2 Training NER models:
+For all experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model and fine-tune it on our synthetic and organic training sets in two distinct settings.
   1. 1.
@@
   2. 2.
-    In the second setting, we first fine-tune the model on the high quality NER data in a language related to the target language [^1] See Appendix for the full list of chosen related languages for all the target languages., so as to get a working NER model in the related language, and then further fine-tune this NER model on our synthetic or organic target language data.
+    In the second setting, we first fine-tune the model on the high quality NER data in a language related source language [^2] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model on our synthetic or organic target language data.
 While the first setting—which we name ner from scratch—aims to shed light on the relative utility of synthetic data for training an NER model (largely) from the ground up, the latter —which we name ner fine-tuning—simulates a common setting, when a lower resource language lacks adequate NER data, but is related to a higher-resource language with existing NER systems.
 In both settings, we modulate the amount of data (both synthetic and organic) used, so as to compare model performance when trained on smaller or larger amounts of each type of data.
-3.2 Implementation
-We source the pre-trained XLM-RoBERTa-large weights from Huggingface using the transformers library; fine-tuning is implemented using training pipelines from the same library.
-In the ner from scratch setting, we train on the the target language data for 10 epochs; in the ner fine-tuning setting, we train on the related language data for 5 epochs, and then the target language data for 10 epochs.
-In all cases, we use a learning rate of 2e-05, and a batch size of 16.
-3.3 Languages & Datasets
+Languages & Datasets:
+We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European).
+Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus.
+We use the Universal NER dataset Mayhew et al.
+(2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al.
+(2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al.
+(2023) for Tamil, Kannada, Malayalam and Telugu.
+While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise.
+Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
+Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
+Additionally, we compare models trained on LLM-generated data with those trained using WikiANN Pan et al.
+(2017); Rahimi et al.
+(2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study.
+This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases.
+Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al.
+(2022); Asai et al.
+(2024).
+4 Results
+4.1 Synthetic Data Generation
+We generate the synthetic datasets following the process described in Section 3.1.
+While model responses from GPT-4 are almost always usable, we found recurring errors in responses from the other two models.
+Some of these errors are described in Table 1 in Appendix A; we discard such instances when compiling our synthetic datasets from model responses.
+The average percentage of usable training datapoints from GPT-4, Llama-3.1 and aya-expanse are 97%, 59.3% and 11.7% respectively.
+[^3] Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits.
+We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
+4.2 Training on Synthetic Data
+Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings.
+While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that models trained on GPT-4-generated data come the closest to models trained on organic data.
+We also find that more synthetic data is not necessarily useful; for some languages, we see a saturation after about 1000 data points, and for some, we also notice a drop in performance with more data.
+Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language outperforms the same models after they have been further fine-tuned on synthetic target language data.
+This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance.
+(a) ner from scratch Setting
+(b) ner fine-tuning Setting
+Figure 2: NER model performance when trained on increasingly large subsets of training data.
+aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as organic or GPT-4-produced data in fine-tuning data size.
+Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model.
+Comparison with WikiAnn:
+In most cases, when dataset size is comparable, training on WikiANN data in the ner from scratch setting yielded NER models that perform considerably worse than those trained on synthetic data from GPT-4.
+For the four Niger-Congo languages, GPT-4 generated data gave superior results even in the ner fine-tuning setting (see Table 3 in Appendix C for the detailed results).
+5 Conclusions and Discussion
+Our results lead us to three main conclusions around the utility of LLM-generated synthetic data for low resource language NER.
+  1. 1.
+    A small amount of carefully annotated data yields better performance than a large amount of synthetic data.
+As is evident in Figure 2, even 100 manually annotated data points can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
+  2. 2.
+    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance.
+  3. 3.
+    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, especially for the more low-resource languages among the ones we studied.
+This echoes the findings by Lignos et al.
+(2022), who arrive at similarly negative findings around the data quality of WikiANN.
+Overall, while showing how synthetic data from LLMs can help train NER models from scratch for low resource languages, our results reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
+6 Limitations
+Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates).
+Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories.
+While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data.
+Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation.
+Our findings should be viewed after taking these aspects into consideration.
+7 Ethics and Broader Impact
+We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work.
+The models built using such artificially created datasets should always be validated with a human-labeled data.
+We did not involve any human participants in this study.
+All the code and generated datasets is provided at this github repository to support reproducible research: https://anonymous.4open.science/r/low-resource-syn-ner-A1C7/.
+Appendix A Synthetic Data Generation
 LLM Response Quality Examples
@@
 Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested.
-We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European).
-Of these, Igbo, Yoruba, and Kinyarwanda are not a part of the 100 languages covered in the pre-trained XLM-Roberta model, but Swahili is a part of the pre-training languages.
-We use the Universal NER dataset Mayhew et al.
-(2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al.
-(2022) for Kinyarwanda, Swahili, Igbo and Yoruba, and the Naamapadam dataset Mhaske et al.
-(2023) for Tamil, Kannada, Malayalam and Telugu.
-It is worth noting that while the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus, in the authors’ admission, contain some amount of noise.
-Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
-Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
-The Universal NER and Naamapadam datasets cover persons, locations and organizations as categories; the MasakhaNER2 data covers these three categories, as well as dates.
-3.4 Synthetic Data Generation Process
-Our synthetic data generation process involves using large language models to generate new synthetic datapoints on the basis of existing, high quality NER annotations through the following steps:
-  1. 1.
-    First, we randomly sample $m$ data points from the train split of an organic NER dataset.
-  2. 2.
-    Next, we format and append these data points to a prompt asking the model to produce $n$ new, unique data points on the basis of the $m$ datapoints in the prompt.
-  3. 3.
-    We submit this prompt as input to the LLM, and extract any correctly-formatted data points from its response;
-  4. 4.
-    We repeat steps (1)-(3) $k$ times, with each call to the model choosing a different random sample of organic data points.
-Figure 1 shows a demonstration of this approach, with the exact prompt used.
-In our experiments, we set $m$ to 10, $n$ to 20, and $k$ to 500.
-This sets an upper cap of 5000 synthetic training data points, if every model response contains perfectly formatted data points.
-Our method to extract correctly-formatted datapoints from raw model responses relies on regular expressions to identify potential JSON substrings in the model response; as shown in Table 1, model responses may contain data that is ill-formatted due to a number of reasons.
-We deem any data faulty for such reasons (among any other reasons, such as hallucinated labels) to be unusable, and do not include it in our synthetic train sets.
-We experiment with three LLMs as our source of synthetic data.
-We use GPT-4222We use gpt-4-turbo, and all data generation with the model was conducted between November and December 2024.
-Achiam et al.
-(2023), which we assume to be the state of the art; we use Llama-3.1-8B-Instruct Dubey et al.
-(2024), as an instance of a much smaller, open-source instruction-tuned model; and finally, we use Cohere For AI’s aya-expanse-32b Dang et al.
-(2024), as an instance of a larger open source multilingual large language.
-4 Results
-4.1 Qualitative Analysis of LLM-Generated Data
-Model Average % of Usable Train Datapoints
-(Across All Languages)
-GPT-4-turbo 97.0%
-Llama-3.1-8B-Instruct* 59.3%
-aya-expanse-32b 11.7%
-Table 2: Average percentage of usable train datapoints produced by each model; we divide the number of usable synthetic datapoints produced for each language by 5000 (the synthetic dataset size if all responses were perfect), and then take the average across languages.
-*Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits.
-(a) ner from scratch Setting
-(b) ner fine-tuning Setting
-Figure 2: NER model performance on a subset of languages when trained on increasingly large subsets of organic and synthetic data.
-aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as the organic or GPT-4-produced data in fine-tuning data size.
-Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model.
-See Appendix for similar plots of all 11 languages.
-Table 1 illustrates the type of responses we get from the three models when we prompt them for synthetic data in the manner described in Section 3.4.
-Broadly speaking, we find that GPT-4 provides the highest quality datapoints, followed by Llama-3.1-8B-Instruct, and lastly aya-expanse-32b.
-This is equally borne out by the proportion of usable training datapoints produced by each model, shown in Table 2.
-While model responses from GPT-4 are almost always usable, we find a few persistent errors in responses from the other two models. We highlight some below:
-  1. 1.
-    Unequal token and tag lengths: Some responses from a model are largely well-formatted and coherent, but are rendered unusable because of a mismatch in their numbers of tokens and tags.
-  2. 2.
-    Run-on responses and incomplete data: Some responses from the models result in endless repeated sequences; in other cases, responses hit new token limits before arriving at complete datapoints.
-  3. 3.
-    Empty responses and prompt continuations: These were specific to aya-expanse-32b; model responses to the generation prompt often either terminated with <EOS> tokens, or simply continued the generation prompt, as shown in the example in Table 1.
-Given the wide range of languages we test, and the cost and difficulty of finding trained annotators to manually annotate samples from all 11 languages, we do not source human annotations of the synthetic data to compare against.
-Instead, as detailed in Section 3.1, we assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
-4.2 Training on Synthetic Data
-Figure 2 shows our results on a subset of languages—Danish, Tamil and Igbo—when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings (see Appendix for such plots across all languages).
-As the figure illustrates, in all cases, models trained on LLM-generated data struggle to compete with those trained on the original, organic train sets.
-Perhaps more surprisingly, in the ner fine-tuning setting, zero-shot transfer onto the target language outperforms the same models after they have been further fine-tuned on synthetic target language data, suggesting that in such cases, the synthetic data is actually detrimental to their performance.
-In terms of comparisons between data-generating LLMs, these results also confirm our initial qualitative observations (see Section 4.1).
-Models trained on synthetic data from GPT-4 do better than those trained on synthetic data from Llama-3.1-8B-Instruct, which in turn outperform the aya-expanse-32b equivalents—suggesting that the quality of data produced by the three LLMs follows the same order.
-4.3 WikiAnn
-Language n.f.s. F1 n.f.t. F1 Data Size
-Kannada WikiANN 4.5e-3 0.77 100
-GPT-4 0.26 0.65 100
-GPT-4 0.65 0.68 4861
-Naamapadam 0.47 0.79 100
-Naamapadam 0.76 0.79 5000
-Danish WikiANN 0.72 0.71 20000
-GPT-4 0.60 0.68 4857
-Universal NER 0.83 0.85 4383
-Yoruba WikiANN 0.07 0.21 100
-GPT-4 0.26 0.43 100
-GPT-4 0.53 0.56 4761
-MasakhaNER 2 0.20 0.50 100
-MasakhaNER 2 0.79 0.82 5000
-Swahili WikiANN 0.50 0.59 1000
-GPT-4 0.74 0.78 1000
-GPT-4 0.75 0.79 4900
-MasakhaNER 2 0.69 0.85 1000
-MasakhaNER 2 0.92 0.90 5000
-Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for a handful of languages.
-See Appendix for full table of results.
-n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.
-As a further evaluation of the viability of synthetically produced data for lower-resource language NER, we also extend our comparison to the WikiANN dataset.
-The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al.
-(2017); Rahimi et al.
-(2019) [^3] As Lignos et al.
-(2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al.
-(2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al.
-(2019), contains data from 176 languages.
-In this work, we refer to the latter when referring to the WikiANN dataset..
-Though it is often used as a multilingual NER benchmark (e.g.
-Asai et al., 2024), however, the WikiANN dataset in fact does not include manual annotations, and is instead produced by scraping Wikipedia data.
-Consequently, the dataset represents a different sort of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases.
-Table 3 shows the performance of NER models trained on WikiANN data, as compared with those trained on our organic and LLM-produced synthetic data.
-The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g.
-Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints).
-In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
-As Table 3 indicates, training on WikiANN data yields NER models that perform considerably worse than those trained on organic data, and often worse than those trained on synthetic data produced by GPT-4 (when data size is equal between the two).
-The WikiANN data for some languages appears completely ill-suited for training an NER model in the ner from scratch setting, while, perhaps unsurprisingly, organic data produces the best results.
-5 Conclusions and Discussion
-Our results suggest three broad findings around the utility of LLM-generated synthetic data for low resource language NER.
-  1. 1.
-    The first, and perhaps most significant, is that a small amount of carefully annotated data yields better performance than a large amount of synthetic data.
-As is evident in Figure 2, even 100 manually annotated datapoints can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
-  2. 2.
-    The second is that, in many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may actually lower its performance.We find this to be true in the case of all languages tested except the four Niger-Congo languages—Igbo, Yoruba, Swahili and Kinyarwanda.
-In these cases, it is worth noting that the overall baselines are lower, presumably because these languages are all lower resource than the others tested.This may explain why synthetic data yields performance gains over the zero-shot baseline, though it does not change the trend of a small amount of manually annotated data yielding far better performance.
-  3. 3.
-    Lastly, we find that LLM-generated synthetic data often produces better results than automatically labelled data.
-Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN data, as shown in the case of Kannada, Yoruba and Swahili.
-This echoes the findings by Lignos et al.
-(2022), who arrive at similarly negative findings around the data quality of WikiANN; consequently, they further reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
-6 Limitations
-Although we experimented with many languages, the nature of the NER datasets used is relatively simple.
-Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories.
-While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data.
-Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation.
-So, our findings should be viewed after taking these aspects into consideration.
-7 Ethics and Broader Impact
-We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work.
-The models built using such artificially created datasets should always be validated with a human-labeled data.
-We did not involve any human participants in this study.
-All the code and generated datasets will be provided as supplementary material to support reproducible research.
-Acknowledgments
-This research was conducted at the National Research Council of Canada, thereby establishing a copyright belonging to the Crown in Right of Canada, that is, to the Government of Canada.
-Appendix A Related-Language Model Details
+As shown in Figure 1, we present the following prompt to the LLM in the data generation process:
+Help me make a {language} Named Entity Recognition dataset.
+Please give me {n} new datapoints, formatted as a single JSON object.
+Make sure the examples are unique and diverse.
+Here are some examples to get you started:
+{m examples}
+For GPT-4, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings.
+For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt.
+For the open-sourced models, we used the following system prompt:
+You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce. You only output JSON strings.
+For GPT-4, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
+You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce.
+We ran all open-sourced models using vLLM Kwon et al.
+(2023), with a temperature setting of 0.8, maximum new token limit of 4096 new tokens, and nucleus sampling value of 0.8.
+Calls to GPT-4 were made using default hyperparameters.
+Table 1 shows some of the examples of the different types of responses to these prompts.
+Appendix B Related-Language Model Details
 In the ner fine-tuning setting, we first train an NER model on a language related to the target language, before fine-tuning it further on the target language NER data.
@@
 *English is not closely related to Slovak, but given the absence of another highly related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
-Appendix B LLM Prompts & Implementation
-As shown in Figure 1, we present the following prompt to the LLM in the data generation process:
-Help me make a {language} Named Entity Recognition dataset.
-Please give me {n} new datapoints, formatted as a single JSON object.
-Make sure the examples are unique and diverse.
-Here are some examples to get you started:
-{m examples}
-For GPT-4, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings.
-For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt.
-For the open-sourced models, we used the following system prompt:
-You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce. You only output JSON strings.
-For GPT-4, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
-You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce.
-We ran all open-sourced models using vLLM Kwon et al.
-(2023), with a temperature setting of 0.8, maximum new token limit of 4096 new tokens, and nucleus sampling value of 0.8.
-Calls to GPT-4 were made using default hyperparameters.
-Appendix C Full Results of Training on Synthetic Data
-C.1 NER From Scratch Setting
-Figure 3 is a plot of results of training on synthetic data, in the ner from scratch setting, for all languages.
-C.2 NER Fine-tuning Setting
-Figure 4 is a plot of results of training on synthetic data, in the ner fine-tuning setting, for all languages.
-Appendix D Full Results of WikiANN Comparison
-Below in Table 5 is the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data.
+B.1 NER-fine tuning: Implementation Details
+We source the pre-trained XLM-RoBERTa-large weights from Huggingface using the transformers library; fine-tuning is implemented using training pipelines from the same library.
+In the ner from scratch setting, we train on the the target language data for 10 epochs; in the ner fine-tuning setting, we train on the related language data for 5 epochs, and then the target language data for 10 epochs.
+In all cases, we use a learning rate of 2e-05, and a batch size of 16.
+Appendix C Full Results of WikiANN Comparison
+The WikiANN dataset is a massively multilingual NER benchmark, comprising data from 176 languages Pan et al.
+(2017); Rahimi et al.
+(2019).
+[^4] As Lignos et al.
+(2022) also note, strictly speaking, the original version of WikiANN put together by Pan et al.
+(2017) contains data from 282 languages; the version of the dataset commonly downloaded from Huggingface, however, and put together by Rahimi et al.
+(2019), contains data from 176 languages.
+In this work, we refer to the latter when referring to the WikiANN dataset.
+Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data.
+The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g.
+Kannada and Yoruba, whose WikiANN train sets contain only 100 datapoints).
+In such cases, however, we compare model performance when trained on equally small amounts of organic or LLM-produced synthetic data.
 Language n.f.s. F1 n.f.t. F1 Data Size
@@
 Telugu WikiANN 0.67 0.74 1000
-GPT-4 0.15 0.68 100
+GPT-4 0.64 0.66 1000
 GPT-4 0.67 0.72 4919
-Naamapadam 0.50 0.79 100
+Naamapadam 0.77 0.82 1000
 Naamapadam 0.83 0.82 5000
@@
 n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.
-Figure 3: ner from scratch Setting
-Figure 4: ner fine-tuning Setting

### v3

@@
 Abstract
-In this paper, we explore whether synthetic datasets generated by large language models are useful for low-resource named entity recognition, considering 11 languages from diverse language families.
-Our results suggest that synthetic data created with seed human labeled data is a reasonable choice when there is no available labeled data, and is better than using automatically labeled data.
+We explore whether synthetic datasets generated by large language models using a few high quality seed samples are useful for low-resource named entity recognition, considering 11 languages from three language families.
+Our results suggest that synthetic data created with such seed data is a reasonable choice when there is no available labeled data, and is better than using entirely automatically labeled data.
 However, a small amount of high-quality data, coupled with cross-lingual transfer from a related language, always offers better performance.
+[^1] Data and code available at: https://github.com/grvkamath/low-resource-syn-ner.
 Does Synthetic Data Help
@@
 1 Introduction
-Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP).
+Named Entity Recognition (NER) for low-resource languages aims to produce robust systems for languages with limited labeled training data available, and has been an area of increasing interest within natural language processing (NLP) over the past decade.
 Two common approaches to address this data scarcity are cross-lingual transfer and data augmentation/synthesis; recent research has in particular explored the usefulness of large language models (LLMs) for such data augmentation and synthetic data creation in NLP Whitehouse et al.
@@
 (2023), while their use for NER is also emerging Bogdanov et al.
+(2024); Dao et al.
 (2025).
+Figure 1: High-level overview of our data generation process.
+We use multilingual large language models to generate new NER data points on the basis of a handful of high quality human labeled data points.
+See Section 3.1 for more.
 In this background, we propose LLM-based synthetic data generation using a small amount of gold examples (Figure 1) as an alternative to relying on automatically created datasets for low-resource NER.
-With experiments covering 11 languages, we show that
+With experiments covering 11 languages from 3 language families—Danish, Swedish and Slovak from the Indo-European language family; Swahili, Kinyarwanda, Yoruba and Igbo from the Niger-Congo language family, and Kannada, Malayalam, Tamil, and Telugu from the Dravidian language family, we show that:
   1. 1.
@@
   3. 3.
-    Synthetic data generated by prompting an LLM with a few high quality examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
-Figure 1: High-level overview of our data generation process.
-We use multilingual large language models to generate new NER datapoints on the basis of a handful of high quality data points.
-See Section 3.1 for more.
+    Synthetic data generated by prompting an LLM with a few high quality (generally human labeled) examples (Figure 1) could be better than using automatically labeled datasets when training low-resource NER models.
 We start with a review of related literature (Section 2) and describe our data generation approach and experimental setup in Section 3, followed by a discussion of the results (Section 4), limitations (Section 6) and broader impact (Section 7).
@@
 (2022).
-More recent work has explored using LLMs such as GPT-3.5 and GPT-4 as data generators for NER Bogdanov et al.
+There are also large existing automatically constructed multilingual NER datasets that rely on sources such as Wikipedia Pan et al.
+(2017); Krishnan et al.
+(2021); Malmasi et al.
+(2022), some of which have become a part of large multilingual benchmarks Asai et al.
+(2024).
+More recent work has explored using LLMs as data generators for NER Bogdanov et al.
 (2024); Heng et al.
+(2024); Evuru et al.
 (2024).
 We build on such work, but differ from their methods.
-Our data generation process uses high quality examples as seeds, and we not only evaluate different LLMs (both open and closed-source), but also experiment with 11 languages covering three language families and five base scripts.
+Our data generation process uses high quality, human validated examples as seeds, and we not only evaluate different LLMs (both open and closed-source) as synthetic data generators, but also experiment with 11 languages covering three language families and five base scripts.
+To our knowledge, this is the first attempt to explore using large language models for synthetic data generation in low-resource NER, and the first to cover >10>10 languages.
 3 Our Approach
@@
   • Next, we format and append these data points to a prompt asking the model to produce $n$ new, unique data points on the basis of the $m$ data points in the prompt.
-  • We submit this prompt as input to the LLM, and extract any correctly-formatted data points from its response;
+  • We submit this prompt as input to the LLM, and extract the correctly-formatted data points from its response;
   • We repeat steps (1)-(3) $k$ times, with each call to the model choosing a different random sample of organic data points.
@@
 Appendix A provides further details about this process.
-We compare three LLMs as our source of synthetic data: GPT-4111We use gpt-4-turbo, and all data generation with the model was conducted between September and December 2024.
-Achiam et al.
-(2023), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al.
-(2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al.
-(2024), as a larger open source multilingual LLM.
-3.2 Training NER models:
-For all experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model and fine-tune it on our synthetic and organic training sets in two distinct settings.
-  1. 1.
-    In the first setting, we use our data to train an NER model from scratch, by fine-tuning the pre-trained XLM-RoBERTa-large on target language NER data.
-  2. 2.
-    In the second setting, we first fine-tune the model on the high quality NER data in a language related source language [^2] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model on our synthetic or organic target language data.
-While the first setting—which we name ner from scratch—aims to shed light on the relative utility of synthetic data for training an NER model (largely) from the ground up, the latter —which we name ner fine-tuning—simulates a common setting, when a lower resource language lacks adequate NER data, but is related to a higher-resource language with existing NER systems.
-In both settings, we modulate the amount of data (both synthetic and organic) used, so as to compare model performance when trained on smaller or larger amounts of each type of data.
-Languages & Datasets:
-We focus on 11 languages from diverse language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European).
-Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus.
-We use the Universal NER dataset Mayhew et al.
-(2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al.
-(2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al.
-(2023) for Tamil, Kannada, Malayalam and Telugu.
-While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise.
-Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
-Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
-Additionally, we compare models trained on LLM-generated data with those trained using WikiANN Pan et al.
-(2017); Rahimi et al.
-(2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study.
-This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases.
-Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al.
-(2022); Asai et al.
-(2024).
-4 Results
-4.1 Synthetic Data Generation
-We generate the synthetic datasets following the process described in Section 3.1.
-While model responses from GPT-4 are almost always usable, we found recurring errors in responses from the other two models.
-Some of these errors are described in Table 1 in Appendix A; we discard such instances when compiling our synthetic datasets from model responses.
-The average percentage of usable training datapoints from GPT-4, Llama-3.1 and aya-expanse are 97%, 59.3% and 11.7% respectively.
-[^3] Llama-3.1’s rejected datapoints are often incomplete due to hitting new token limits, suggesting potentially higher capabilities under higher token limits.
-We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
-4.2 Training on Synthetic Data
-Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings.
-While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that models trained on GPT-4-generated data come the closest to models trained on organic data.
-We also find that more synthetic data is not necessarily useful; for some languages, we see a saturation after about 1000 data points, and for some, we also notice a drop in performance with more data.
-Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language outperforms the same models after they have been further fine-tuned on synthetic target language data.
-This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance.
-(a) ner from scratch Setting
-(b) ner fine-tuning Setting
-Figure 2: NER model performance when trained on increasingly large subsets of training data.
-aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do not extend as far as organic or GPT-4-produced data in fine-tuning data size.
-Performance at Fine-tuning Dataset Size = 0, only present in the ner fine-tuning setting, indicates zero-shot performance of a related-language NER model.
-Comparison with WikiAnn:
-In most cases, when dataset size is comparable, training on WikiANN data in the ner from scratch setting yielded NER models that perform considerably worse than those trained on synthetic data from GPT-4.
-For the four Niger-Congo languages, GPT-4 generated data gave superior results even in the ner fine-tuning setting (see Table 3 in Appendix C for the detailed results).
-5 Conclusions and Discussion
-Our results lead us to three main conclusions around the utility of LLM-generated synthetic data for low resource language NER.
-  1. 1.
-    A small amount of carefully annotated data yields better performance than a large amount of synthetic data.
-As is evident in Figure 2, even 100 manually annotated data points can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
-  2. 2.
-    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance.
-  3. 3.
-    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, especially for the more low-resource languages among the ones we studied.
-This echoes the findings by Lignos et al.
-(2022), who arrive at similarly negative findings around the data quality of WikiANN.
-Overall, while showing how synthetic data from LLMs can help train NER models from scratch for low resource languages, our results reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
-6 Limitations
-Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates).
-Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories.
-While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data.
-Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation.
-Our findings should be viewed after taking these aspects into consideration.
-7 Ethics and Broader Impact
-We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work.
-The models built using such artificially created datasets should always be validated with a human-labeled data.
-We did not involve any human participants in this study.
-All the code and generated datasets is provided at this github repository to support reproducible research: https://anonymous.4open.science/r/low-resource-syn-ner-A1C7/.
-Appendix A Synthetic Data Generation
 LLM Response Quality Examples
@@
 <EOS_TOKEN>include a mix of names, locations, organizations...
-Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested.
+Table 1: Examples of different types of responses from the synthetic data-generating LLMs tested, across languages
+We compare three LLMs as our source of synthetic data: GPT-4.1222We use gpt-4.1-2025-04-14.
+Note that in an earlier draft of this work, we used gpt-4-turbo Achiam et al.
+(2023), when it represented the state-of-the-art; surprisingly, gpt-4-turbo yielded slightly better results.
+Nevertheless, here we report results on GPT-4.1, to better represent currently available models.
+OpenAI (2025), which we assume to be the state of the art; Llama-3.1-8B-Instruct Dubey et al.
+(2024), as a much smaller, open-source instruction-tuned model; and finally, aya-expanse-32b Dang et al.
+(2024), as a larger open source multilingual LLM.
+3.2 Training NER models:
+For all experiments, we use the pre-trained version of XLM-RoBERTa-large Conneau et al. (2020) as our base model and fine-tune it on our synthetic and organic training sets in two distinct settings.
+  1. 1.
+    In the first setting, we use our synthetic data to train an NER model from scratch, by fine-tuning XLM-RoBERTa-large on target language NER data.
+  2. 2.
+    In the second setting, we first fine-tune the model on the high quality NER data in a related source language [^3] See Table 2 in Appendix B for the full list of chosen related languages for all the target languages., and then further fine-tune this NER model using our synthetic or organic target language data.
+While the first setting—which we name ner from scratch—aims to shed light on the relative utility of synthetic data for training an NER model (largely) from the ground up, the latter —which we name ner fine-tuning—simulates a common setting, when a lower resource language lacks adequate NER data, but is related to a higher-resource language with existing NER systems.
+In both settings, we modulate the amount of data (both synthetic and organic) used, so as to compare model performance when trained on smaller or larger amounts of each type of data.
+Languages & Datasets:
+We focus on 11 languages from three distinct language families: Tamil, Kannada, Malayalam, Telugu (Dravidian), Kinyarwanda, Swahili, Igbo, Yoruba (Niger-Congo), Swedish, Danish and Slovak (Indo-European).
+Of these, Igbo, Yoruba, and Kinyarwanda are not among the 100 languages in the XLM-Roberta pre-training corpus.
+We use the Universal NER dataset Mayhew et al.
+(2024) as our high quality, manually annotated dataset for Swedish, Danish and Slovak; MasakhaNER2 Adelani et al.
+(2022) for Kinyarwanda, Swahili, Igbo and Yoruba; and the Naamapadam dataset Mhaske et al.
+(2023) for Tamil, Kannada, Malayalam and Telugu.
+While the first two datasets are completely manually annotated, the train and validation splits of the Naamapadam dataset are constructed using parallel corpora, and thus contain some noise.
+Nevertheless, we choose it as our organic dataset, as (i) its test sets, which contain 500-1000 datapoints per language, are completely manually annotated, and (ii) it remains the largest NER resource for these four languages.
+Crucially, all of these datasets cover largely identical NER categories, allowing for comparisons between them.
+The Universal NER and Naamapadam datasets cover persons, locations and organizations as categories; the MasakhaNER2 data covers these three categories, as well as dates.
+Additionally, we compare models trained entirely on LLM-generated data with those trained using WikiANN Pan et al.
+(2017); Rahimi et al.
+(2019), a large, automatically created NER dataset based on Wikipedia cross-linking, as it covers the 11 languages we study.
+This dataset represents a different form of synthetic data—one generated not from LLMs, but instead from scraping knowledge bases without any seed data.
+Although the dataset has no manual annotations, it is frequently used as a standard low-resource NER benchmark Schmidt et al.
+(2022); Asai et al.
+(2024).
+4 Results
+(a) ner from scratch Setting
+(b) ner fine-tuning Setting
+Figure 2: NER model performance when trained on increasingly large subsets of training data.
+aya-expanse-32b and Llama-3.1-8B-Instruct produced lower amounts of usable data; this is why they do often not extend as far as organic or GPT-4.1-produced data in fine-tuning data size.
+In the ner fine-tuning setting, performance at Fine-tuning Training Data Size = 0 indicates zero-shot performance of a related-language NER model.
+4.1 Synthetic Data Generation
+We generate the synthetic datasets following the process described in Section 3.1.
+While model responses from GPT-4.1 are generally usable, we found more recurring errors in responses from the other two models.
+Some of these errors are described in Table 1; we discard such instances when compiling our synthetic datasets from model responses.
+The average percentage of usable training datapoints from GPT-4.1, Llama-3.1 and aya-expanse are 82.6%, 59.7% and 11.1% respectively.
+We assess the overall quality and viability of this synthetic data by measuring the performance of an NER model on a high quality, manually-annotated test set, when trained on the synthetic data.
+4.2 Training on Synthetic Data
+Figure 2 shows our results when using synthetic data from different models, in both the ner from scratch and ner fine-tuning settings.
+While the models trained on organic data in the ner from scratch setting always perform better than synthetic data based models, we find that the models trained on GPT-4.1-generated data often come the closest to models trained on organic data compared to the other synthetic data sources.
+Best results with synthetic data based training are seen for Danish, followed by Swahili and Swedish.
+We also find that more synthetic data is not necessarily useful; for most languages, we see relative saturation after 1000 data points, and in the case of Kannada, we actually notice a drop in performance with more data.
+Perhaps more surprisingly, in the ner fine-tuning setting, we notice that zero-shot transfer from a related language usually outperforms the same models after they have been further fine-tuned on synthetic target language data.
+Fine-tuning the related language NER model with organic data from the target language helped for only Kinyarwanda and Yoruba.
+This suggests that in some cases where an NER model for a related language exists, synthetic data in target languages may actually be detrimental to overall performance.
+Models trained on synthetic data from GPT-4.1 do better than those trained on synthetic data from Llama-3.1-8B-Instruct only about half of the time; on the other hand, there are often too few usable aya-expanse-32b datapoints for a fair comparison.
+Comparison with WikiAnn:
+In addition to comparing models fine-tuned on synthetic data versus organic data, we also looked into the question of whether our synthetic data generation approach offers any benefits over automatically labeled datasets, taking WikiAnn as an example.
+Training models on WikiANN data leads to higher performance than training on GPT-4.1-generated data only for the four Dravidian languages in our data, but generally leads to significantly worse performance than training on synthetic data from GPT-4.1 for the remaining languages (see Table 3 in Appendix C for detailed results).
+This holds in both the ner from scratch and ner fine-tuning settings, when data size is comparable; and, in the case of Danish and Swedish, training on WikiANN leads to worse performance even when there is several times more WikiANN data than GPT-4.1-generated data.
+Overall, we can conclude that using synthetic data following our approach appears to be better than relying on WikiAnn for most languages.
+This echoes the findings by Lignos et al.
+(2022), who arrive at similarly negative findings around the data quality of WikiANN, and calls for not considering results on WikiANN as a benchmark for multilingual NER comparisons in the future.
+5 Conclusions and Discussion
+Our results lead us to three main conclusions around the utility of LLM-generated synthetic data for low resource language NER.
+  1. 1.
+    A small amount of carefully annotated data yields better performance than a large amount of synthetic data.
+As is evident in Figure 2, even 100 manually annotated data points can yield NER models that cannot be matched by models trained on much larger amounts of synthetic data.
+  2. 2.
+    In many cases, zero-shot transfer from a related-language NER model is a high baseline, and that further training such a model on synthetic data may even lower the performance.
+We find this to be true in the case of all languages tested except the Yoruba and Swahili.
+For these two languages, it is worth noting that the overall baselines are lower, presumably because these languages are all lower resource than the others tested.
+This may explain why synthetic data yields performance gains over the zero-shot baseline, though it does not change the trend of a small amount of manually annotated data yielding far better performance.
+  3. 3.
+    Despite the fact that it falls short of manually annotated data, LLM-generated data often still yields better model performance than WikiANN, which is automatically extracted from Wikipedia texts.
+Overall, while showing how synthetic data from LLMs can help train NER models from scratch for low resource languages, our results reinforce the need for manually annotated gold test sets in benchmarking NER for lower resource languages.
+6 Limitations
+Although we experimented with many languages, the nature of the NER datasets used is relatively simple, containing only three or four entity categories (persons, locations, organizations and dates).
+Thus, we don’t know if the general conclusions, especially about the quality of synthetic data, will extend to scenarios where there are many entity categories.
+While we did study datasets covering more than one language family, the selection of language is far from extensive, and is also constrained by the availability of human labeled test data.
+The observations need not necessarily hold true across all language families, naturally.
+Finally, to keep the experiments under control, we explored a limited set of methods for fine-tuning and synthetic data generation.
+Our findings should be viewed after taking these aspects into consideration.
+7 Ethics and Broader Impact
+We used publicly available datasets with human-annotated and automatically labeled data, and also created synthetically generated datasets as a part of this work.
+The models built using such artificially created datasets should always be validated with a human-labeled data.
+We did not involve any human participants in this study.
+All the code and generated datasets is provided at this GitHub repository to support reproducible research: https://github.com/grvkamath/low-resource-syn-ner.
+Acknowledgments
+This research was conducted at the National Research Council of Canada, thereby establishing a copyright belonging to the Crown in Right of Canada, that is, to the Government of Canada.
+Gaurav Kamath is supported by a Doctoral Training Award from the Fonds de Recherche du Québec—Société et Culture.
+Appendix A Synthetic Data Generation
 As shown in Figure 1, we present the following prompt to the LLM in the data generation process:
@@
 {m examples}
-For GPT-4, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings.
+We prompt GPT-4.1 using OpenAI’s batch API functionality [^4] https://platform.openai.com/docs/guides/batch; for the open-source models, we use the vLLM library Kwon et al. (2023) to run inference.
+For GPT-4.1, we used the OpenAI API’s functionalities for structured outputs to ensure that outputs were formatted as JSON strings.
 For the open-sourced models, we experimented with using transformers-compatible libraries for obtaining structured outputs from LLMs, but ultimately found better results simply specifying the JSON requirement in the model and system prompt.
@@
 You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce. You only output JSON strings.
-For GPT-4, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
+For GPT-4.1, we used the following (minimally different) system prompt, on the assumption that specifying output mode in the system prompt was less important on account of the API’s structured output functionalities:
 You are a helpful model that helps build text-based datasets, but does not produce any conversation besides the text it is asked to produce.
-We ran all open-sourced models using vLLM Kwon et al.
-(2023), with a temperature setting of 0.8, maximum new token limit of 4096 new tokens, and nucleus sampling value of 0.8.
-Calls to GPT-4 were made using default hyperparameters.
+We ran both open-sourced models with a temperature setting of 0.8, and nucleus sampling value of 0.8.
+We initially used a maximum new token limit of 4096 for both models.
+However, noticing that some of Llama-3.1-8B-Instruct’s unusable datapoints were specifically due to hitting new token limits, we regenerated data from this model with a maximum new token limit of 8192.
+Calls to GPT-4.1 were made using default hyperparameters.
 Table 1 shows some of the examples of the different types of responses to these prompts.
@@
 Table 2: List of related languages used in the ner fine-tuning setting for each target language.
-*English is not closely related to Slovak, but given the absence of another highly related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
+*English is not closely related to Slovak, but given the absence of another closely related language among the 11 target languages, it was chosen as the language for the base NER model to be fine-tuned.
 B.1 NER-fine tuning: Implementation Details
@@
 In this work, we refer to the latter when referring to the WikiANN dataset.
-Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4-produced data, and WikiANN data.
+Table 3 shows the full list of comparisons between NER model performance when trained on organic data, GPT-4.1-produced data, and WikiANN data.
 The sizes of the WikiANN train sets vary significantly between different languages, meaning we often cannot assess the quality of the data in the context of training sets containing over 1000 datapoints (e.g.
@@
 Kannada WikiANN 4.5e-3 0.77 100
-GPT-4 0.26 0.65 100
-GPT-4 0.65 0.68 4861
-Naamapadam 0.47 0.79 100
-Naamapadam 0.76 0.79 5000
+GPT-4.1 0.03 0.75 100
+GPT-4.1 0.00 0.43 4899
+Naamapadam 0.49 0.79 100
+Naamapadam 0.80 0.78 5000
 Telugu WikiANN 0.67 0.74 1000
-GPT-4 0.64 0.66 1000
-GPT-4 0.67 0.72 4919
-Naamapadam 0.77 0.82 1000
-Naamapadam 0.83 0.82 5000
+GPT-4.1 0.00 0.40 1000
+GPT-4.1 0.18 0.47 4931
+Naamapadam 0.71 0.81 1000
+Naamapadam 0.84 0.83 5000
 Tamil WikiANN 0.55 0.62 15000
-GPT-4 0.56 0.51 4977
-Naamapadam 0.73 0.73 5000
+GPT-4.1 0.00 0.57 4944
+Naamapadam 0.73 0.74 5000
 Malayalam WikiANN 0.65 0.74 10000
-GPT-4 0.64 0.70 4898
-Naamapadam 0.79 0.83 5000
+GPT-4.1 0.25 0.35 4859
+Naamapadam 0.83 0.83 5000
 Yoruba WikiANN 0.07 0.21 100
-GPT-4 0.26 0.43 100
-GPT-4 0.53 0.56 4761
-MasakhaNER 2 0.20 0.50 100
-MasakhaNER 2 0.79 0.82 5000
+GPT-4.1 0.20 0.42 100
+GPT-4.1 0.37 0.50 4187
+MasakhaNER 2 0.28 0.54 100
+MasakhaNER 2 0.83 0.84 5000
 Swahili WikiANN 0.50 0.59 1000
-GPT-4 0.74 0.78 1000
-GPT-4 0.75 0.79 4900
-MasakhaNER 2 0.69 0.85 1000
-MasakhaNER 2 0.92 0.90 5000
+GPT-4.1 0.74 0.78 1000
+GPT-4.1 0.78 0.79 2619
+MasakhaNER 2 0.88 0.91 1000
+MasakhaNER 2 0.92 0.91 5000
 Kinyarwanda WikiANN 7.9e-4 0.35 100
-GPT-4 0.23 0.46 100
-GPT-4 0.58 0.54 4754
-MasakhaNER 2 0.26 0.61 100
-MasakhaNER 2 0.80 0.81 5000
+GPT-4.1 0.29 0.46 100
+GPT-4.1 0.47 0.50 2683
+MasakhaNER 2 0.07 0.61 100
+MasakhaNER 2 0.79 0.81 5000
 Igbo WikiANN 7.7e-3 0.39 100
-GPT-4 0.43 0.70 100
-GPT-4 0.66 0.71 4693
-MasakhaNER 2 0.41 0.72 100
-MasakhaNER 2 0.81 0.86 5000
+GPT-4.1 0.29 0.55 100
+GPT-4.1 0.53 0.55 3473
+MasakhaNER 2 0.48 0.76 100
+MasakhaNER 2 0.82 0.83 5000
 Danish WikiANN 0.72 0.71 20000
-GPT-4 0.60 0.68 4857
+GPT-4.1 0.79 0.82 4112
 Universal NER 0.83 0.85 4383
 Swedish WikiANN 0.36 0.29 20000
-GPT-4 0.65 0.56 4825
-Universal NER 0.58 0.89 4303
+GPT-4.1 0.69 0.74 4215
+Universal NER 0.87 0.90 4303
 Slovak WikiANN 0.57 0.55 20000
-GPT-4 0.29 0.29 4889
-Universal NER 0.80 0.82 5000
-Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4, and high quality ‘organic’ data, for all 11 languages.
+GPT-4.1 0.29 0.29 4508
+Universal NER 0.77 0.80 5000
+Table 3: Performance of NER models trained on WikiANN, synthetic data from GPT-4.1, and high quality ‘organic’ data, for all 11 languages.
 n.f.s: ner from scratch setting; n.f.t: ner fine-tuning setting.